OpenStudy (anonymous):

What is the mass of an elevator moving upwards with an acceleration of 2 meters per second, which experiences a tension of 6000N by the elevator cable. Please explain. Thanks!

OpenStudy (anonymous):

Use formula \[\Sigma F=ma\] rearrange the formula since you need to find the mass not the force. so it should look like this \[m=\Sigma F \div a\] Then you just plug in the numbers in your calculator :D The answer should be 3000kg

OpenStudy (anonymous):

there are two forces

OpenStudy (anonymous):

F-mg=ma F= (a+g)m m=F/(a+g) m=6000/(2+10)=500kg

OpenStudy (anonymous):

dodomaroc got the right answer :D

OpenStudy (anonymous):

oh yeah, can you explain the variables used?